OpenStudy (anonymous):

Solve the system of equations using substitution. 3y = –x + 2 y = –x + 9

Parth (parthkohli):

It's given that \(y = -x + 9\), so that means \(3y = -x + 2 \implies 3(-x + 9) = -x + 2 \)

OpenStudy (anonymous):

That is not an option for any of my answers

Parth (parthkohli):

You need to solve it; don't look directly into the options.

OpenStudy (anonymous):

I dont know how to solve it. That is why I posted the question. ;)